OpenStudy (anonymous):

"What mass of PbCl2 contains 1.0X10^24 atoms Cl?"

OpenStudy (aaronq):

first, find the mass percent compositions of Cl in PbCl. then find how many grams 1.0*10^24 atoms of Cl equate to. the rest should be straight forward

OpenStudy (aaronq):

composition* not plural

OpenStudy (anonymous):

Huh? I am sorry, but I am still confused. I know the answer should be 230g. Mass Percent of Cl in PbCl2= 26%? What do I do?

OpenStudy (aaronq):

whats the mass of that many Cl atoms?

OpenStudy (anonymous):

71

OpenStudy (anonymous):

or i am not sure :(

OpenStudy (aaronq):

\(n=\dfrac{N}{N_A}=\dfrac{1.0*10^{24}}{6.023*10^{23}}\)=1.66 moles N= number of atoms (or molecules), \(N_A\) = avogadros number \(m_{Cl}=n_{Cl}*M_{Cl}\)=1.66 moles*35.45g/mol=58.847 g

OpenStudy (aaronq):

good so far?

OpenStudy (anonymous):

somewhat yes

OpenStudy (anonymous):

how to go from that to eventually 230g?

OpenStudy (anonymous):

and what does m stand for in the second equation

OpenStudy (aaronq):

m = mass (in g), M=molar mass

OpenStudy (aaronq):

so, 26%*m=58.847 g m=58.847g/0.26=226.3346153846153846

OpenStudy (aaronq):

i didn't round, so my answer is a bit off from what you expected (230 g)

OpenStudy (anonymous):

thanks, i'm going to try to write it out on paper to see if i understand :). Ill be right back

OpenStudy (aaronq):

okay !

OpenStudy (anonymous):

so I haven't learned those two equations (in the first part), where did you get them from and when are we supposed to use it?

OpenStudy (anonymous):

and the 1.66 moles is moles of Cl? and not PbCl2?

OpenStudy (aaronq):

these equations? \(n=\dfrac{N}{N_A};\;\; m=n*M\)

OpenStudy (aaronq):

yeah 1.66 moles of Cl because you're using the amount of atoms of Cl not the whole molecule

OpenStudy (anonymous):

ya! those equations. My teacher hasn't showed us that yet. and why did you calcualte 26% (mass percent) + why did divide w/ the mass percent? Also, why didn't we use the molar mass of PbCl2 and just stayed w/ mass of Cl2 the whole time? thanks!

OpenStudy (aaronq):

we'll the first one isn't really an equation, it's just using the definition of a mole. 1 mole = 6.023*10^23 atoms, so it's merely a conversion. They should be in your book, it's one of the first concepts you see in stoichiometry. the 2nd is just a rearrangement of \(n=\dfrac{m}{M}\), also seen early in stoichiometry. We only used molar mass of Cl because were asked about Cl atoms not PbCl2 molecules. We used the mass percent because we needed to lol. for example, i tell you that 50% by mass of my backpack is books and i tell you that i have 10 books (atoms) each weighing (1 g). Whats the total mass of my backpack? 10 books = 10g 10g=50% of the backpack so to find the mass of the backpack 10g=\(m_{backpack}\)50% \(\rightarrow m_{backpack}=\dfrac{10g}{50\%}\ \) 50%=0.5 \(m_{backpack}=\dfrac{10g}{0.5}\)

OpenStudy (anonymous):

simple! use unitary (ratio method) i.e: if (6.02x10^23 x 2 ) ( avagadros number times no of Cl atoms) atoms are there in 278.2 g (Mr) of PbCl2 then 1.0x10^24 woud be there in : (278.2) * (1.0x10^24)/(6.02x10^23) put that in calculator and get the ans! hope this helps!